Let X be a standard normal random variable. Find E[2X2 + 1] UESTION 2 Roll two fair dice. If we denote random variables E, F, and G as E = { Sum of two numbers is 6), F- { First die shows 4), and G Sum of two numbers is 7) OE and F are independent OE and F are not independent O E and F are mutually exclusive Unable to tell with the given information UESTION 3 Which of the following statements is true? O F and G are independent O F and G are not independent O F and G are mutually exclusive Unable to tell with the given information
Answer #1

1)

Solution :

X be the standard normal variable.

E(X 2) = 1

E[2X 2 + 1] = 2E(X 2) + E(1) = 2 * 1 + 1 = 2 + 1 = 3